OpenStudy (anonymous):

y=1/2*5 to the power of x

OpenStudy (anonymous):

need more info.

jhonyy9 (jhonyy9):

if 5 is exponent you can writing y=1/2*5 =2*(-5) - so to the power x will be y=1/2*(x) =2*(-x)